# Learning Dense Hand Contact Estimation from Imbalanced Data

### v1

@@ A.7 Computational requirements
Table A3 reports the computational requirements of HACO with various backbone configurations. The original HACO model, listed in the first row, shows the highest computational cost, particularly in training memory, which exceeds 26GB. This is primarily due to the memory overhead associated with storing gradients while fine-tuning the ViT-H dosovitskiy2020image backbone pretrained on HaMeR pavlakos2024reconstructing . To support broader adoption in downstream tasks such as 3D hand grasp generation grady2021contactopt and hand-object reconstruction liu2024easyhoi , we provide additional HACO variants with reduced computational demand. We first evaluate the FPN lin2017feature backbone from HandOccNet park2022handoccnet . Although FPN is known to be computationally expensive due to its multi-scale feature hierarchy, it remains significantly more efficient than ViT-H. We also test lighter ViT variants such as ViT-B and ViT-S, which offer favorable trade-offs by maintaining moderate memory usage while achieving inference speeds exceeding 60 fps. We further evaluate HRNet sun2019deep backbones, including HRNet-W48 and HRNet-W32. Despite their moderate memory consumption, both models exhibit low inference speeds, falling below 20 fps. Finally, we include ResNet-based he2016deep variants ranging from lightweight models such as ResNet-18 to deeper configurations such as ResNet-152. These models generally show low memory usage, fewer parameters, and high runtime efficiency. In particular, ResNet-18 achieves an inference speed exceeding 100 fps while maintaining a computational cost under 3 GFLOPs.
@@
Model Backbone Train Memory (MB) Model Memory (MB) Params. (M) Speed (fps) GFLOPs
HACO ViT-H xu2022vitpose 26,157 2,629 671.10 54.26 125.63
HACO FPN lin2017feature 6,887 230 59.56 66.71 9.94
HACO ViT-L xu2022vitpose 16,649 1,305 341.91 65.79 62.87
HACO ViT-B xu2022vitpose 7,767 469 122.84 96.80 18.54
HACO ViT-S xu2022vitpose 4,960 218 56.34 86.88 5.10
HACO HRNet-W48 sun2019deep 11,296 468 122.37 10.52 23.48
HACO HRNet-W32 sun2019deep 9,026 334 86.13 17.14 12.55
HACO ResNet-152 he2016deep 9,252 395 103.05 29.84 15.93
HACO ResNet-101 he2016deep 7,471 337 87.40 41.14 11.07
HACO ResNet-50 he2016deep 5,919 264 68.41 73.25 6.21
HACO ResNet-34 he2016deep 4,163 219 56.75 80.85 5.03
HACO ResNet-18 he2016deep 3,785 178 46.64 111.23 2.61
A.8 More qualitative results
@@
Additional qualitative comparisons of dense hand contact estimation between HACO and prior methods, including POSA hassan2021populating , BSTRO huang2022capturing , and DECO tripathi2023deco , are shown in Figures A4, A5, and A6, which correspond to the MOW cao2021reconstructing , Hi4D yin2023hi4d , and both HIC tzionas2016capturing and RICH huang2022capturing datasets, respectively. Overall, HACO consistently outperforms existing approaches by a substantial margin. Unlike prior methods that frequently fail to predict contact even when the hand is clearly interacting, HACO reliably captures accurate and dense hand contact. Among the baselines, DECO demonstrates stronger performance than POSA and BSTRO, likely as a result of being trained on in-the-wild datasets with more diverse contact distributions. However, it remains to be affected by class imbalance, resulting in a high rate of false negatives. BSTRO is more severely affected by this imbalance, frequently predicting no contact across the entire hand and producing overly coarse contact regions concentrated in the palmar area. This behavior likely stems from its reliance on human-scene interaction datasets during training. POSA also exhibits frequent false negatives and fails to detect contact even in visibly interacting regions, likely due to its dependence on a pose prior rather than direct contact supervision with visual input.
Figure A4: Qualitative comparison of dense hand contact estimation with POSA hassan2021populating , BSTRO huang2022capturing , and DECO tripathi2023deco on MOW cao2021reconstructing dataset. We highlight exemplar regions where HACO outperform previous methods. Note that we only predict right hand contact.
@@
Figure A5: Qualitative comparison of dense hand contact estimation with POSA hassan2021populating , BSTRO huang2022capturing , and DECO tripathi2023deco on Hi4D yin2023hi4d dataset. We highlight exemplar regions where HACO outperform previous methods. Note that we only predict right hand contact.
@@
Figure A6: Qualitative comparison of dense hand contact estimation with POSA hassan2021populating , BSTRO huang2022capturing , and DECO tripathi2023deco on HIC tzionas2016capturing and RICH huang2022capturing dataset. We highlight exemplar regions where HACO outperform previous methods. Note that we only predict right hand contact.
A.9 Limitations and societal impacts
Limitations. Our HACO supports dense hand contact estimation across a wide range of interaction types, including hand-object, hand-hand, hand-face, hand-scene, and hand-body interactions. However, hands frequently engage in self-contact under certain poses (e.g., contact between the thumb and index finger in an "okay" gesture) or during whole-body motions muller2021self . We intentionally exclude self-contact cases, as including them could negatively impact downstream applications. For example, ContactOpt grady2021contactopt may incorrectly optimize grasps between the hand and an object if self-contact is misinterpreted as external contact. As a result, HACO currently does not support dense hand contact estimation in self-contact scenarios. Nevertheless, we consider modeling self-contact a promising direction for future research, particularly for applications in the metaverse (e.g., AR/VR) and action recognition.

### v2

@@ A.9 Computational requirements
Table A7 reports the computational requirements of HACO with various backbone configurations. The original HACO model, listed in the first row, shows the highest computational cost, particularly in training memory, which exceeds 26GB. This is primarily due to the memory overhead associated with storing gradients while fine-tuning the ViT-H [10] backbone pretrained on HaMeR [49]. To support broader adoption in downstream tasks such as 3D hand grasp generation [14] and 3D hand-object reconstruction [35], we provide additional HACO variants with reduced computational demand. We first evaluate the FPN [31] backbone from HandOccNet [46]. Although FPN is known to be computationally expensive due to its multi-scale feature hierarchy, it remains significantly more efficient than ViT-H. We also test lighter ViT variants such as ViT-B and ViT-S, which offer favorable trade-offs by maintaining moderate memory usage while achieving inference speeds exceeding 60 fps. We further evaluate HRNet [57] backbones, including HRNet-W48 and HRNet-W32. Despite their moderate memory consumption, both models exhibit low inference speeds, falling below 20 fps. Finally, we include ResNet-based [20] variants ranging from lightweight models such as ResNet-18 to deeper configurations such as ResNet-152. These models generally show low memory usage, fewer parameters, and high runtime efficiency. In particular, ResNet-18 achieves an inference speed exceeding 100 fps while maintaining a computational cost under 3 GFLOPs.
@@
Model Backbone Train Memory (MB) Test Memory (MB) Params. (M) Speed (fps) GFLOPs
HACO ViT-H [63] 26,157 12,728 671.10 54.26 125.63
HACO FPN [31] 6,887 3,002 59.56 66.71 9.94
HACO ViT-L [63] 16,649 4,604 341.91 65.79 62.87
HACO ViT-B [63] 7,767 2,940 122.84 96.80 18.54
HACO ViT-S [63] 4,960 2,434 56.34 86.88 5.10
HACO HRNet-W48 [57] 11,296 7,574 122.37 10.52 23.48
HACO HRNet-W32 [57] 9,026 7,806 86.13 17.14 12.55
HACO ResNet-152 [20] 9,252 5,096 103.05 29.84 15.93
HACO ResNet-101 [20] 7,471 4,978 87.40 41.14 11.07
HACO ResNet-50 [20] 5,919 4,826 68.41 73.25 6.21
HACO ResNet-34 [20] 4,163 2,456 56.75 80.85 5.03
HACO ResNet-18 [20] 3,785 2,372 46.64 111.23 2.61
A.10 More qualitative results
@@
Additional qualitative comparisons of dense hand contact estimation between HACO and prior methods, including POSA [18], BSTRO [22], and DECO [58], are shown in Figures A4, A5, and A6, which correspond to the MOW [4], Hi4D [65], and both HIC [59] and RICH [22] datasets, respectively. Overall, HACO consistently outperforms existing approaches by a substantial margin. Unlike prior methods that frequently fail to predict contact even when the hand is clearly interacting, HACO reliably captures accurate dense hand contact. Among the baselines, DECO demonstrates stronger performance than POSA and BSTRO, likely as a result of being trained on in-the-wild datasets with more diverse contact distributions. However, it remains to be affected by class imbalance, resulting in a high rate of false negatives. BSTRO is more severely affected by this imbalance, frequently predicting no contact across the entire hand and producing overly coarse contact regions concentrated in the palmar area. This behavior likely stems from its reliance on human-scene interaction datasets during training. POSA also exhibits frequent false negatives and fails to detect contact even in visibly interacting regions, likely due to its dependence on a pose prior rather than direct contact supervision with visual input.
@@
Figure A4: Qualitative comparison of dense hand contact estimation with POSA [18], BSTRO [22], DECO [58] on MOW [4] dataset. We highlight exemplar regions where HACO outperforms previous methods. Note that we only predict right hand contact.
@@
Figure A5: Qualitative comparison of dense hand contact estimation with POSA [18], BSTRO [22], DECO [58] on Hi4D [65] dataset. We highlight exemplar regions where HACO outperforms previous methods. Note that we only predict right hand contact.
Figure A6: Qualitative comparison of dense hand contact estimation with POSA [18], BSTRO [22], DECO [58] on HIC [59] and RICH [22] dataset. We highlight exemplar regions where HACO outperforms previous methods. Note that we only predict right hand contact.
A.11 Limitations and societal impacts
Limitations. Our HACO supports dense hand contact estimation across a wide range of interaction types, including hand-object, hand-hand, hand-face, hand-scene, and hand-body interactions. However, hands frequently engage in self-contact under certain poses (e.g., contact between the thumb and index finger in an "okay" gesture) or during whole-body motions [42]. We intentionally exclude self-contact cases, as including them could negatively impact downstream applications. For example, ContactOpt [14] may incorrectly optimize grasps between the hand and an object if self-contact is misinterpreted as external contact. As a result, HACO currently does not support dense hand contact estimation in self-contact scenarios. Nevertheless, we consider modeling self-contact a promising direction for future research, particularly for applications in the metaverse (e.g., AR/VR) and action recognition. Furthermore, temporal context is essential to learn accurate dense hand contact estimation as contact is most noticeable when the movement of hand is obstructed by an object or scene. Integration of temporal-based methods [67, 44] that facilitate temporal information from input video can stabilize the performance of dense hand contact estimation. With the advent of diffusion-based methods [25, 39], designing a diffusion-based approach for dense hand contact estimation is another important research direction as it enables the learning of dense hand contact priors from the large-scale data. Additionally, our balanced contact sampling (BCS) utilizes a single reference point of dataset-wide dense hand contact mean. While this allows HACO to fairly represent diverse contact statistics, if distribution of the statistics is multi-modal, this strategy might not be effective. There can be a room for improvement on our BCS strategy if the multi-modality of dense hand contact datasets is proven and multi-modal aware methodology can be incorporated to BCS. Moreover, we evaluate HACO only on samples that contain at least one vertex in contact. However, this evaluation protocol may overlook the model’s inability to handle false hand-level contact predictions, which are important to consider for real-world applications. The incorporation of additional evaluation protocols such as accuracy, specificity, and Matthews Correlation Coefficient may be beneficial. Lastly, dense hand contact estimation demonstrates inferior performance on interacting hand scenarios relative to other hand interaction scenarios. The incorporation of 3D interacting hand mesh reconstruction method [11, 45] or the use of additional training dataset for hand-hand interaction [40] could enhance the performance.